Please help me find the focus and the dirctrix of the parabola 16x^2=y.
There are 4 choices for an answer:
1. Focus (0,4), directrix y=-4
2. Focus (0,1/4), directrix y=-1/4
3. Focus (0,1/16), directrix y=-1/16
4. Focus (0,1/64), directrix y=-1/64
Thank you for your time and help, Natalie!
: Please help me find the focus and the dirctrix of the parabola 16x^2=y.
There are 4 choices for an answer:
1. Focus (0,4), directrix y=-4
2. Focus (0,1/4), directrix y=-1/4
3. Focus (0,1/16), directrix y=-1/16
4. Focus (0,1/64), directrix y=-1/64
Thank you for your time and help, Natalie!

Answer by Alan3354(1218)

16x^2=y.
There are 4 choices for an answer:
1. Focus (0,4), directrix y=-4
2. Focus (0,1/4), directrix y=-1/4
3. Focus (0,1/16), directrix y=-1/16
4. Focus (0,1/64), directrix y=-1/64
Thank you for your time and help, Natalie!
----------------
The standard from for parabolas is x^2=4py
So y = x^2/4p = 16x^2
x^2/4p = 16x^2
4p = x^2/16x^2 = 1/16
p = 1/64
P is the distance from the Origin to the directrix, and (0,p) is the focus.
That's answer #4.
Solved by pluggable solver: SOLVE quadratic equation (work shown, graph etc)
Quadratic equation ax^2+bx+c=0 (in our case 16x^2+0x+0 = 0) has the following solutons:

x[12] = (b+-sqrt( b^2-4ac ))/2\a

For these solutions to exist, the discriminant b^2-4ac should not be a negative number.

First, we need to compute the discriminant b^2-4ac: b^2-4ac=(0)^2-4*16*0=0.

Discriminant d=0 is zero! That means that there is only one solution: x = (-(0))/2\16.
Expression can be factored: 16x^2+0x+0 = (x-0)*(x-0)

Again, the answer is: 0, 0. Here's your graph:
graph( 500, 500, -10, 10, -20, 20, 16*x^2+0*x+0 )